# Adversarial bandit optimization for approximately linear functions

### v5

@@ 4.2 Analysis
$Regret=\underbrace{\sum_{t=1}^{T}(x^{\prime}_{t}-h)^{\top}g_{t}}_{\textsc{Reg-Term}}+\underbrace{\sum_{t=1}^{T}(y^{\prime}_{t}-x^{\prime}_{t})^{\top}\theta^{\prime}_{t}}_{\textsc{Deviation-Term}}+\underbrace{\sum_{t=1}^{T}(d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t})^{\top}(x^{\prime}_{t}-h)}_{\textsc{Error-Term}},$ (3)
where $h\in\mathcal{K^{\prime}}$, $\theta^{\prime}=(\theta,0)$. This differs from the decomposition proposed by Lee et al. [17], which divides the regret as follows:
@@
Secondly, by obtaining the expected bound and high-probability bound for Deviation-Term, we can derive the expected regret bound and high-probability regret bound, respectively. For the high-probability bound of Deviation-Term, unlike SCRiBLe with lifting and increasing learning rates [17], which constrains the decision set from $\mathcal{K^{\prime}}$ to $\mathcal{K^{\prime}_{\delta}}$ to ensure that $x^{\prime}_{t}$ is never too close to the boundary (thus ensuring that the eigenvalues of $\mathbf{A}_{t}$ are bounded, especially for bounding $\|h\|_{\nabla^{2}\mathcal{R}(x^{\prime}_{t})}$). Our approach does not require $x^{\prime}_{t}$ to stay away from the boundary. Furthermore, we do not need to bound the eigenvalues of $\mathbf{A}_{t}$, which gives us greater flexibility in choosing the value of $\delta$ (such as $\frac{1}{T^{2}}$), leading to a better upper bound for the regret.
@@ 5.1 $\nu$-normal barrier
Lemma 1 ([19, 20])
@@
Lemma 3[13])
Let $\mathcal{R}$ is a $\nu$-normal barrier over $\Psi$, then for all $x,z\in int(\Psi):\mathcal{R}(z)-\mathcal{R}(x)\leq\nu\log\frac{1}{1-\pi_{x}(z)}$, where $\pi_{x}(z)=\inf\{t\geq 0:x+t^{-1}(z-x)\in\Psi\}$.
$\nu$-normal barrier plays a crucial role in addressing one of the key challenges in this problem: bounding $\|x_{t}^{\prime}-h\|_{x_{t}^{\prime}}$, where $h\in\mathcal{K^{\prime}}$. Equation 8 give $\|x_{t}^{\prime}\|_{x_{t}^{\prime}}=\sqrt{\nu}$. Building upon Lemma 1, we introduce an effective Lemma 4 that aids in bounding $\|h\|_{x_{t}^{\prime}}$.
Lemma 4
If $\mathcal{R}$ be a $\nu$-normal barrier for $\Psi\subseteq R^{d}$, and $x\in int(\Psi)$, $h\in\Psi$ satisfy $\nabla\mathcal{R}(x)^{\top}(h-x)\geq-\nu$, then
$\|h\|_{x}\leq 2\nu.$ (13)
With the help of Equation 8 and Lemma 4, it is easy to apply the triangle inequality to derive $\|x_{t}^{\prime}-h\|_{x_{t}^{\prime}}\leq\|x_{t}^{\prime}\|_{x_{t}^{\prime}}+\|h\|_{x_{t}^{\prime}}\leq\sqrt{\nu}+2\nu$, where $h\in\mathcal{K^{\prime}}$.
@@ 5.2 Useful lemmas
Like Lemma 6 in the SCRiBLe algorithm [2], the next minimizer $x^{\prime}_{t+1}$ is “close” to $x^{\prime}_{t}$ and Lemma 5 implies $\|x^{\prime}_{t+1}-x^{\prime}_{t}\|_{x^{\prime}_{t}}\leq 4d\eta$. This result will help us bound $g_{t}^{\top}(x^{\prime}_{t}-h)$, where $h\in\mathcal{K^{\prime}}$(see Lemma 7).
@@
This next lemma is based on Lemma B.9. [17], but due to the differences in the loss functions, $g_{t,i}$ is no longer an unbiased estimate of $\theta_{t,i}$, for $i\in[d]$. Lee et al. state that $\mathbb{E}_{t}[l_{t,i}]=\theta_{t,i}$, for $i\in[d]$, where $l_{t}=d(\theta_{t},0)(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}$ [17]. We directly apply it to our analysis.
@@
The regret bound of FTRL algorithm [13] states that for every $u\in\mathcal{K}$, $\sum_{t=1}^{T}\nabla_{t}^{\top}x_{t}-\sum_{t=1}^{T}\nabla_{t}^{\top}u\leq\sum_{t=1}^{T}[\nabla_{t}^{\top}x_{t}-\nabla_{t}^{\top}x_{t+1}]+\frac{1}{\eta}[\mathcal{R}(u)-\mathcal{R}(x_{1})]$, where $\nabla_{t}$ represents the gradient of the loss function $f_{t}$. In our adaptation, we replaced $\nabla_{t}$ with $g_{t}$ and $\mathcal{K}$ with $\mathcal{K^{\prime}}$. This modification does not fundamentally alter the original result. Since the update way $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$ satisfied the condition of FTRL algorithm [13], we can apply Lemma 5.3. in  [13] to Algorithm 1 as follow.
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}\leq 2\eta d^{2}T+\frac{\nu\log(\frac{1}{\delta})}{\eta}+Td\epsilon(2\nu+\sqrt{\nu})+\delta DGT.$ (14)
@@
$\mathbb{E}[\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}]$$\leq$ $\mathbb{E}[4\eta d^{2}T+\frac{\nu\log(\frac{1}{\delta})}{\eta}+Td\epsilon(2\nu+\sqrt{\nu})+\delta DGT]$
$\leq$ $4\eta d^{2}T+\frac{\nu\log(\frac{1}{\delta})}{\eta}+Td\epsilon(2\nu+\sqrt{\nu})+\delta DGT.$
@@
$\leq$ $4d\sqrt{2\nu T\log T}+\frac{GD}{T}+Td\epsilon(2\nu+\sqrt{\nu})+2T\epsilon,$
where $\eta=\frac{\sqrt{2\nu\log T}}{2d\sqrt{T}},\delta=\frac{1}{T^{2}}$.
@@
$\sum_{t=1}^{T}X_{t}\leq C(\sqrt{8Vln(C/\gamma)}+2B^{*}ln(C/\gamma)),$ (16)
where $V=\max\{1;\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}\mid F_{t}]\},B^{*}=\max_{t\in[T]}B_{t}$, and $C=\lceil logb\rceil\lceil log(b^{2}T)\rceil$.
The analysis in [17] employs Lemma 9 to derive a high-probability bound for $\sum_{t=1}^{T}[y_{t}^{\top}\theta_{t}-x_{t}^{\top}g_{t}+u^{\top}(g_{t}-\theta_{t})]$. In contrast, our approach defines $X_{t}=\theta_{t}^{\top}y_{t}-\theta_{t}^{\top}x_{t}$ and derives the high-probability bound for $\sum_{t=1}^{T}(\theta_{t}^{\top}y_{t}-\theta_{t}^{\top}x_{t})$. This distinction in the application of Lemma 9 enables us to derive a tighter high-probability upper bound for bandit linear optimization.
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}\leq 4d\sqrt{2\nu T\ln T}+\frac{GD}{T}+Td\epsilon(2\nu+\sqrt{\nu}),$ (17)
where $\eta=\frac{\sqrt{2\nu\ln T}}{2d\sqrt{T}},\delta=\frac{1}{T^{2}}$. Then by Lemma 9,
@@ 5.5 Application to black-box optimization
Let $\hat{x}$ be the output of Algorithm 1. Then, it follows from Theorem 1 that $f(\hat{x})-\min_{x\in\mathcal{K}}f(x)\leq\frac{4d\sqrt{2\nu\ln T}}{\sqrt{T}}+\frac{GD}{T^{2}}+d\epsilon(2\nu+\sqrt{\nu})+2\epsilon$. Additionally, we provide a lower bound $2\epsilon$ for this problem (see Lemma 10). We can see that the difference between the lower bound and the upper bound is only $d\epsilon(2\nu+\sqrt{\nu})$ as $T$ approaches infinity. This suggests the potential existence of "easier" settings between the adversarial environment and the standard stochastic environment, where better algorithms might be found. It also motivates us to explore these settings further.
@@ 6 Lower bound
Firstly, suppose that the algorithm $\mathcal{A}$ is deterministic. At iteration $t=1,...,T$, for any feedback $y_{1},...,y_{t-1}\in\mathbb{R}$, $\mathcal{A}$ should choose the next query point $x_{t}$ based on the data observed so far. That is,
@@ 7 Experiments
Figure 1: Average cumulative loss of algorithms for artificial data sets. The blue line corresponds to the results of our algorithm, the yellow line represents the algorithm in [17], and the green line corresponds to the SCRiBLe algorithm [2].
We conduct comparative experiments across a range of $\epsilon$ values: 0, 0.25, 0.5, 0.75, and 1, evaluating Algorithm 1, SCRiBLe [2], and SCRiBLe with lifting and increasing learning rates [17]. We generate the artificial data as follows. The input set is constructed such that each element $x\in\mathbb{R}^{d}$ satisfies $\|x\|_{2}<D$. The loss vectors $\theta_{1},...,\theta_{T}$ are $d$-dimensional vectors randomly generated before the experiment, each with a Euclidean norm no greater than $G$. The $\nu$-normal barrier is defined as $\mathcal{R}(x,b)=400\cdot\left(-\log\left(1-\frac{\|x\|^{2}}{D^{2}}\right)-2\nu\log b\right)$ and $\eta$ is defined as $\frac{20\sqrt{2\log T}}{4d\sqrt{T}}$. In our experiments, we set $D=3$, $G=1$, $d=5$, respectively. The perturbation is defined as $\sigma(y_{t})=\epsilon\sin{((y_{t}^{\top}l)\pi)}$, where $l$ is a random vector in $\mathbb{R}^{d}$ (Note that this is not the worst-case perturbation, which may explain why the results do not increase significantly as $\epsilon$ grows). Each algorithm is run with a fixed time horizon of $T=2000$. For each value of $\epsilon$, we repeat the experiment 10 times for each algorithm and compute the average cumulative loss. The x-axis represents the value of $\epsilon$, and the y-axis shows the average cumulative loss. As shown in the figure above, our method generally outperforms SCRiBLe. Compared to Lee et al.’s method, our approach achieves slightly better performance. This indicates that, when facing oblivious adversaries, not using increasing learning rates (i.e., dynamically adjusting the value of $\eta$) has little to no impact on the results(see Fig. 1).
@@ 0.A.1 Proof of Lemma 4
From Lemma 1, we have
$\|h\|_{x}\leq-\nabla\mathcal{R}(x)^{\top}h\leq\lvert\nabla\mathcal{R}(x)^{\top}h\rvert.$ (23)
Then,
$\lvert\nabla\mathcal{R}(x)^{\top}h\rvert\\ =\lvert\nabla\mathcal{R}(x)^{\top}(h-x+x)\rvert\\ \leq\lvert\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\lvert\nabla\mathcal{R}(x)^{\top}x\rvert.$ (24)
By Lemma 1, $\lvert\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\lvert\nabla\mathcal{R}(x)^{\top}x\rvert\leq\nu+\lvert x^{\top}\nabla^{2}\mathcal{R}(x)x\rvert=2\nu$.

### v8

@@ 4.2 Analysis
$Regret=\underbrace{\sum_{t=1}^{T}(x^{\prime}_{t}-h)^{\top}E[g_{t}]}_{\textsc{Reg-Term}}+\underbrace{\sum_{t=1}^{T}(y^{\prime}_{t}-x^{\prime}_{t})^{\top}\theta^{\prime}_{t}}_{\textsc{Deviation-Term}}+\underbrace{\sum_{t=1}^{T}(d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t})^{\top}(x^{\prime}_{t}-h)}_{\textsc{Error-Term}},$ (3)
where $h\in\mathcal{K^{\prime}}$, $\theta^{\prime}=(\theta,0)$. This differs from the decomposition proposed by Lee et al. [lee2020bias], which divides the regret as follows:
@@
Secondly, by obtaining the expected bound and high-probability bound for Deviation-Term, we can derive the expected regret bound and high-probability regret bound, respectively. For the high-probability bound of the Deviation-Term when $\epsilon=0$, our approach differs from SCRiBLe with lifting and increasing learning rates [lee2020bias], which constrains $\delta$ to satisfy certain conditions to ensure that $x^{\prime}_{t}$ is never too close to the boundary (thus ensuring that the eigenvalues of $\mathbf{A}_{t}$ are bounded, especially for bounding $\|h\|_{\nabla^{2}\mathcal{R}(x^{\prime}_{t})}$). Our approach does not require $x^{\prime}_{t}$ to stay away from the boundary. Furthermore, we do not need to bound the eigenvalues of $\mathbf{A}_{t}$, which gives us greater flexibility in choosing the value of $\delta$ (such as $\frac{1}{T^{2}}$), leading to a better upper bound for the regret.
@@ 5.1 $\nu$-normal barrier
Lemma 1([nemirovski2004interior, nesterov1994interior])
@@
Lemma 3[hazan2016introduction])
Let $\mathcal{R}$ is a $\nu$-normal barrier over $\Psi$, then for all $x,z\in int(\Psi):\mathcal{R}(z)-\mathcal{R}(x)\leq\nu\ln\frac{1}{1-\pi_{x}(z)}$, where $\pi_{x}(z)=\inf\{t\geq 0:x+t^{-1}(z-x)\in\Psi\}$.
@@ 5.2 Useful lemmas
Lemma 4 plays a crucial role in our analysis, especially in the proof of Lemma 8.
Lemma 4
For Algorithm 1 and for any $x,y\in\mathcal{K^{\prime}_{\delta}}$
$\|x-y\|_{x}\leq 2(\frac{1}{\delta}-1)\left(\nu+2\sqrt{\nu}\right).$ (13)
Like Lemma 6 in the SCRiBLe algorithm [abernethy2008competing], the next minimizer $x^{\prime}_{t+1}$ is “close” to $x^{\prime}_{t}$ and Lemma 5 implies $\|x^{\prime}_{t+1}-x^{\prime}_{t}\|_{x^{\prime}_{t}}\leq 4d\eta$. This result will help us bound $g_{t}^{\top}(x^{\prime}_{t}-h)$, where $h\in\mathcal{K^{\prime}}$(see Lemma 7).
@@
This next lemma is based on Lemma B.9. [lee2020bias], but due to the differences in the loss functions, $g_{t,i}$ is no longer an unbiased estimate of $\theta_{t,i}$, for $i\in[d]$. Lee et al. state that $\mathbb{E}_{t}[l_{t,i}]=\theta_{t,i}$, for $i\in[d]$, where $l_{t}=d(\theta_{t},0)(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}$ [lee2020bias]. We directly apply it to our analysis.
@@
The regret bound of FTRL algorithm [hazan2016introduction] states that for every $u\in\mathcal{K}$, $\sum_{t=1}^{T}\nabla_{t}^{\top}x_{t}-\sum_{t=1}^{T}\nabla_{t}^{\top}u\leq\sum_{t=1}^{T}[\nabla_{t}^{\top}x_{t}-\nabla_{t}^{\top}x_{t+1}]+\frac{1}{\eta}[\mathcal{R}(u)-\mathcal{R}(x_{1})]$, where $\nabla_{t}$ represents the gradient of the loss function $f_{t}$. In our adaptation, we replaced $\nabla_{t}$ with $g_{t}$ and $\mathcal{K}$ with $\mathcal{K^{\prime}}$. This modification does not fundamentally alter the original result. Since the update way $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$ satisfied the condition of FTRL algorithm [hazan2016introduction], we can apply Lemma 5.3. in  [hazan2016introduction] to Algorithm 1 as follow.
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}\leq 4\eta d^{2}T+\frac{\nu\ln(\frac{1}{\delta})}{\eta}+2dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta DGT.$ (14)
@@
$\mathbb{E}[\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}]$ $\leq$ $\mathbb{E}[4\eta d^{2}T+\frac{\nu\ln(\frac{1}{\delta})}{\eta}+dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta DGT]$
$\leq$ $4\eta d^{2}T+\frac{\nu\ln(\frac{1}{\delta})}{\eta}+dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta DGT.$
@@
$\leq$ $4d\sqrt{\nu T\ln\frac{1}{\delta}}+2dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta GDT+2T\epsilon,$
where $\eta=\frac{\sqrt{\nu\ln\frac{1}{\delta}}}{2d\sqrt{T}}$.
@@
$\sum_{t=1}^{T}X_{t}\leq C(\sqrt{8V\ln(C/\gamma)}+2B^{*}\ln(C/\gamma)),$ (16)
where $V=\max\{1;\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}\mid F_{t}]\},B^{*}=\max_{t\in[T]}B_{t}$, and $C=\lceil\log b\rceil\lceil\log(b^{2}T)\rceil$.
The analysis in [lee2020bias] employs Lemma 9 to derive a high-probability bound for $\sum_{t=1}^{T}[y_{t}^{\top}\theta_{t}-x_{t}^{\top}g_{t}+u^{\top}(g_{t}-\theta_{t})]$. In contrast, our approach defines $X_{t}=\theta_{t}^{\top}y_{t}-\theta_{t}^{\top}x_{t}$ and derives the high-probability bound for $\sum_{t=1}^{T}(\theta_{t}^{\top}y_{t}-\theta_{t}^{\top}x_{t})$. This distinction in the application of Lemma 9 enables us to derive a tighter high-probability upper bound for bandit linear optimization.
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}\leq 4\eta d^{2}T+\frac{\nu\ln(\frac{1}{\delta})}{\eta}+2dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta DGT,$ (17)
where $\eta=\frac{\sqrt{\nu\ln T}}{2d\sqrt{T}}$. Then by Lemma 9,
@@ 5.5 Application to black-box optimization
Let $\hat{x}$ be the output of Algorithm 1. Then, it follows from Theorem 1 that $f(\hat{x})-\min_{x\in\mathcal{K}}f(x)\leq\frac{4d\sqrt{\nu\ln\frac{1}{\delta}}}{\sqrt{T}}+\delta GD+d(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+2\epsilon$. Additionally, we provide a lower bound $2\epsilon$ for this problem (see Lemma 10). We can see that the difference between the lower bound and the upper bound is $\delta GD+d(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon$ as $T$ approaches infinity. This suggests the potential existence of "easier" settings between the adversarial environment and the standard stochastic environment, where better algorithms might be found. It also motivates us to explore these settings further.
@@ 6 Lower bound
Firstly, suppose that the algorithm $\mathcal{A}$ is deterministic. At iteration t=1,,Tt=1,...,T, for any feedback $y_{1},...,y_{t-1}\in\mathbb{R}$, $\mathcal{A}$ should choose the next query point $x_{t}$ based on the data observed so far. That is,
@@ 7 Experiments
Figure 1: Average cumulative loss of algorithms for artificial data sets. The blue line corresponds to the results of our algorithm, the yellow line represents the algorithm in [lee2020bias], and the green line corresponds to the SCRiBLe algorithm [abernethy2008competing].
We conduct comparative experiments across a range of $\epsilon$ values: 0, 0.25, 0.5, 0.75, and 1, evaluating Algorithm 1, SCRiBLe [abernethy2008competing], and SCRiBLe with lifting and increasing learning rates [lee2020bias]. We generate the artificial data as follows. The input set is constructed such that each element $x\in\mathbb{R}^{d}$ satisfies $\|x\|_{2}<D$. The loss vectors $\theta_{1},...,\theta_{T}$ are $d$-dimensional vectors randomly generated before the experiment, each with a Euclidean norm no greater than $G$. The $\nu$-normal barrier is defined as $\mathcal{R}(x,b)=400\cdot\left(-\log\left(1-\frac{\|x\|^{2}}{b^{2}D^{2}}\right)-2\nu\log b\right)$ and $\eta$ is defined as $\frac{20\sqrt{\log\frac{1}{\delta}}}{4d\sqrt{T}}$. In our experiments, we set $D=5$, $G=1$, $d=5$, $\nu=1$, respectively. The perturbation is defined as $\sigma(y_{t})=\epsilon\sin{((y_{t}^{\top}l)\pi)}$, where $l$ is a random vector in $\mathbb{R}^{d}$ (Note that this is not the worst-case perturbation, which may explain why the results do not increase significantly as $\epsilon$ grows). Each algorithm is run with a fixed time horizon of $T=2000$. For each value of $\epsilon$, we repeat the experiment 10 times for each algorithm and compute the average cumulative loss. The x-axis represents the value of $\epsilon$, and the y-axis shows the average cumulative loss. As shown in the figure above, our method generally outperforms SCRiBLe. Compared to Lee et al.’s method, our approach achieves slightly better performance. This indicates that, when facing oblivious adversaries, not using increasing learning rates (i.e., dynamically adjusting the value of $\eta$) has little to no impact on the results(see Fig. 1).
@@ Appendix 0.A Appendix
Before prove the Lemma 1, we introduce the follow Lemma first.
Lemma 11([nemirovski2004interior])
For $x\in int(\mathcal{K})$ and $h\in\mathbb{R}^{d}$ let $p_{x}(h)=inf\{r\geq 0|x\pm r^{-1}h\in\mathcal{K}\}$, One has $p_{x}(h)<=|h|_{x}\leq(\nu+2\sqrt{\nu})p_{x}(h)$.
@@ 0.A.1 Proof of Lemma 4
From Lemma 11, for $\forall x,y\in\mathcal{K}_{\delta}$, we have
$\|x-y\|_{x}\leq(\nu+2\sqrt{\nu})p_{x}(x-y).$ (23)
For $p_{x}(x-y)=inf\{r\geq 0|x\pm r^{-1}(x-y)\in\mathcal{K}\}$, it suffices to show $\|x\pm r^{-1}(x-y)\|\leq D$.
$\|x\pm r^{-1}(x-y)\|\leq\|x\|+\|r^{-1}(x-y)\|\leq(1-\delta)D+r^{-1}2(1-\delta)D$ (24)
Then, $r=2\frac{(1-\delta)}{\delta}$ and $\|x-y\|_{x}\leq 2(\nu+2\sqrt{\nu})\frac{(1-\delta)}{\delta}$